# MonoForce: Learnable Image-conditioned Physics Engine

### v3

@@ Abstract
We propose a novel model for the prediction of robot trajectories on rough offroad terrain from the onboard camera images. This model enforces the laws of classical mechanics through a physics-aware neural symbolic layer while preserving the ability to learn from large-scale data as it is end-to-end differentiable. The proposed hybrid model integrates a black-box component that predicts robot-terrain interaction forces with a neural-symbolic layer. This layer includes a differentiable physics engine that computes the robot’s trajectory by querying these forces at the points of contact with the terrain. As the proposed architecture comprises substantial geometrical and physics priors, the resulting model can also be seen as a learnable physics engine conditioned on real images that delivers $10^{4}$ trajectories per second. We argue and empirically demonstrate that this architecture reduces the sim-to-real gap and mitigates out-of-distribution sensitivity. The differentiability, in conjunction with the rapid simulation speed, makes the model well-suited for various applications including model predictive control, trajectory shooting, supervised and reinforcement learning, or SLAM. The codes and data are publicly available [^1] https://github.com/ctu-vras/monoforce.
@@
Robot-terrain Interaction, Differentiable Physics, End-to-end Learning, Monocular Vision, Trajectory Prediction
@@ I Introduction
Autonomous robotics in off-road environments holds immense promise for various commercial applications, including outdoor logistics, inspections, and forestry operations. Yet, unlike systems designed for controlled environments, such as factories, off-road autonomous systems still remain prohibitively immature for dependable deployment. The main challenge that prevents reliable off-road deployment lies in the inability to predict the behavior accurately of the robot on the terrain. For example, the prediction of robot trajectory in tall vegetation or muddy and rocky terrains, especially from camera images and lidar scans, remains an unresolved problem in outdoor robotics.
Figure 1: Qualitative results: The first three rows show (i) onboard camera images (the only input), (ii) predicted supporting terrain and (iii) a 3D view with predicted trajectory and visualized ground truth lidar point clouds (not used for estimation). While the majority of rigid objects are correctly reconstructed from the camera image, robot flippers, tall grass, fallen branches, or soft hanging branches are correctly suppressed as they do not influence the resulting robot trajectory. Predicted terrain friction, dampening, and stiffness are not visualized for brevity.
@@
Over the last decade, roboticists proposed a wide variety of white-box [1, 2, 3] and black-box models [4, 5, 6, 7, 8] for the off-road trajectory prediction. Black-box models primarily suffer from a severe out-of-distribution problem – the phenomenon where the distribution of training data does not correspond to the testing data. This problem naturally arises from the fact that the training data includes only the trajectories from safe terrain traversals, such as small terrain steps, while the decision about terrain traversability also naturally comprises the non-traversable terrains, such as tall cliffs, Figure 4. Since the step-to-cliff-generalization of large state-of-the-art black-box models is typically poor, the safety of the resulting system is arbitrarily inferior. On the other hand, white-box architectures offer good generalization due to substantial inductive bias but often suffer from the sim-to-real gap – the phenomenon where a model """"trained in"""" or """"replaced by"""" simulation faces challenges or discrepancies when applied in the real world. Although several techniques, such as rapid motor adaptation [9], are known to partially suppress this issue through domain randomization, the sim-to-real gap, as well as the out-of-distribution problem, makes trajectory prediction from camera images prohibitively unreliable. We introduce a grey-box model that combines the best of both worlds to achieve better generalization and a smaller sim-to-real gap. This model enforces white-box laws of classical mechanics through a physics-aware neural symbolic layer, while preserving the ability to learn from huge data as it is end-to-end differentiable. We demonstrate that even just a single onboard camera is enough for reliable autonomous deployment in off-road environments; see Figures 13 for some qualitative results and Figure 5 for quantitative evaluation.
We highlight that all state-of-the-art models, including the proposed one, are inherently statistically inconsistent as it is impossible to provide training data covering dangerous cases. Consequently, providing a close-to-infinite number of training examples, as typical, for large foundation models [10] does not imply a superior performance. Addressing the statistical inconsistency for the trajectory prediction task without resorting to real-world robot-damaging trajectories remains an open challenge. In this work, we make a step toward this direction by introducing a grey-box model that leverages: (i) the strengths of black-box models to learn effectively from real data, thereby reducing the sim-to-real gap, and (ii) the strengths of white-box approaches to generalize effectively to unseen data, thereby addressing the out-of-distribution problem.
In our implementation, the black-box model predicts robot-terrain interaction forces and the shape of the robot-supporting terrain, while the neural symbolic layer, which contains a differentiable physics engine, solves the robot’s trajectory by querying these forces at the robot-terrain contacts. The main advantage of this approach is that it enables further self-supervision induced by lidar measurements, which essentially provides an upper bound on the terrain shape. Since lidar measurements are not restricted to safe terrains, the prediction of the shape remains statistically consistent. If we could further assume that terrain textures on unsafe terrains also appear on safe terrains, the whole procedure would be statistically consistent. However, this is not true in general, as there could be some inherently unsafe textures, such as holes covered by vegetation, which do not have their safe counterpart.
@@
As the proposed architecture comprises substantial geometrical and physics priors, the resulting model can be also seen as learnable physics engine conditioned by a real image that delivers $10^{4}$ trajectories per second; see Figure 2 for details (the engine efficiency is also investigated in the Section IV-D). In addition to that, the model is end-to-end-differentiable; therefore, gradients can be backpropagated towards its (i) convolutional filters, (ii) camera and robot parameters, (iii) control, (iv) positions, and (v) terrain properties. The differentiability, in conjunction with the rapid simulation speed, makes the model suitable for a myriad of tasks, including model predictive control [11], trajectory shooting [12], supervised and reinforcement learning [13], SLAM [14], online robot-model reidentification or camera recalibration [15]. The explainable structure of the proposed architecture also delivers a variety of intermediate outputs, such as terrain shape and its physical properties, robot-terrain reaction forces or contact points, which can all serve as efficient sources of self-supervision if measured during the training set creation or restricted in PINN-like manner [16]. We observed that the instability of gradient computation in existing differentiable simulators, such as NVIDIA’s WARP [17] and Google’s BRAX [18], makes them prohibitively unreliable and slow for both the learning and the inference. As a suitable replacement, we introduce a differentiable neural symbolic layer, which involves our from-scratch implemented physics engine, into the proposed architecture. This solution outperforms existing works in gradient stability and computational speed while preserving sufficient accuracy for reliable trajectory prediction. As all target applications, including learning, control, planning, and SLAM, can be naturally parallelized, we also achieve significant speed-up through massive parallelization on GPU.
@@
Step towards statistically consistent learning: Explainability of the proposed grey-box model provides several well-interpretable intermediate outputs that serve as a natural source of self-supervision. The self-supervision, hand in hand with a substantial physical prior, increases the statistical consistency of the learning procedure.
Image-conditioned simulation: The end-to-end differentiable image-conditioned simulation is suitable for a myriad of robotics tasks such as control or SLAM, as it predicts $10^{4}$ trajectories in parallel per second.
Experimental evaluation on non-rigid terrains: The proposed model outperforms other state-of-the-art methods on non-rigid terrains, such as grass or soft undergrowth that deforms when traversed by the robot. See Figure 5 for quantitative results.
@@ II Related Work
[…]
II-A Black-box (Data-driven) Models
The environment representation modeling from a single [23] or multiple RGB images [20] in a top-down view has been widely studied in the literature. The geometry of the scene prediction (taking into account occlusions) from stereo or depth-camera input has been addressed in [24]. The importance and benefits of sensor fusion (camera images, lidar, and radar) for environment representation have been demonstrated in [25]. The authors of RoadRunner [26] developed an end-to-end learning-based framework that predicts the traversability and elevation maps from multiple images and a lidar voxel map. However, the robot-terrain interaction and trajectory prediction tasks are not considered in these works.
The Recurrent Neural Networks (RNN) models [27] and especially their variant Long Short-Term Memory (LSTM) [28] and Gated Recurrent Unit (GRU) [29] have been widely and successfully used [30, 31, 19] for trajectory prediction tasks of various agents (mobile robots, cars, airplanes, people, and more). However, their main drawback lies in the difficulty of capturing the underlying physical laws, which leads to limited generalization capabilities (more in Section III-C and Figure 4).
[…]
Figure 7: Detailed view of MonoForce’s heightmaps: Geometrical heightmap replicates lidar heightmap. Soft heightmap outlines soft parts of the terrain (those that do not generate any robot-terrain interaction forces) such as vegetation. Terrain heightmap consists of two layers: (i) Supporting terrain: geometrical heightmap with subtracted soft parts of the terrain; (ii) Friction heightmap: see significantly higher values on dirt.
@@ II-C Grey-box (Hybrid) Models
This paper substantially extends the preceding work [40] by:
  • integration of other sensor modalities, such as lidar (Section IV-B), into the MonoForce model. It was shown that including the lidar data improves the accuracy of the estimation of terrain properties by 15% (Section IV-B).
  • improvement of the robot model: we use a precise robot model for more accurate contact modeling and physical simulation. The new model is built from the robot’s URDF and uses 223 robot body points in comparison to only 16 points used in the approximate robot model from [40].
  • extension of the predicted terrain properties (friction). It is demonstrated that this improvement, together with the integrated precise robot body model, decreases the trajectory prediction error by 14% in translation and 26% in orientation (Table I).
  • benchmarking against a purely data-driven baseline. It is shown (Section IV-C) that the proposed physics engine outperforms the LSTM-based baseline up to 56% in translation and 48% in orientation in terms of trajectory prediction accuracy, Figure 5.
III Method
A detailed overview of the proposed architecture that converts images and control commands into trajectories is depicted in Figure 6. The model consists of several learnable modules that deeply interact with each other. The terrain encoder carefully transforms visual features from the input image into the heightmap space using known camera geometry. The resultant heightmap features are further refined into interpretable physical quantities that capture properties of the terrain such as its shape, friction, stiffness, and damping. Next, the physics engine combines the terrain properties with the robot model, robot state, and control commands and delivers reaction forces at points of robot-terrain contacts. It then solves the equations of motion dynamics by integrating these forces and delivers the trajectory of the robot. Since the complete computational graph of the feedforward pass is retained, the backpropagation from an arbitrary loss, constructed on top of delivered trajectories, or any other intermediate outputs is at hand.
@@ III-A Terrain Encoder
The part of the MonoForce architecture (Figure 6) that predicts terrain properties $\mathbf{m}$ from sensor measurements $\mathbf{z}$ is called terrain encoder. The proposed architecture starts by converting pixels from a 2D image plane into a heightmap with visual features. Since the camera is calibrated, there is a substantial geometrical prior that connects heightmap cells with the pixels. We incorporate the geometry through the Lift-Splat-Shoot architecture [20]. This architecture uses known camera intrinsic parameters to estimate rays corresponding to particular pixels – pixel rays, Figure 13. For each pixel ray, the convolutional network then predicts depth probabilities and visual features. Visual features are vertically projected on a virtual heightmap for all possible depths along the corresponding ray. The depth-weighted sum of visual features over each heightmap cell is computed, and the resulting multichannel array is further refined by deep convolutional network to estimate the terrain properties $\mathbf{m}$.
The terrain properties include the geometrical heightmap $\mathcal{H}_{g}$, the heights of the terrain supporting layer hidden under the vegetation $\mathcal{H}_{t}=\mathcal{H}_{g}-\Delta\mathcal{H}$, terrain friction $\mathcal{M}$, stiffness $\mathcal{K}$, and dampening $\mathcal{D}$. The intuition behind the introduction of the $\Delta\mathcal{H}$ term is that $\mathcal{H}_{t}$ models a partially flexible layer of terrain (e.g. mud) that is hidden under flexible vegetation, Figure 7.
III-B Differentiable Physics Engine
The differentiable physics engine solves the robot motion equation and estimates the trajectory corresponding to the delivered forces. The trajectory is defined as a sequence of robot states $\tau=\{s_{0},s_{1},\ldots,s_{T}\}$, where $\mathbf{s}_{t}=[\mathbf{x}_{t},\mathbf{v}_{t},R_{t},\bm{\omega}_{t}]$ is the robot state at time $t$, $\mathbf{x}_{t}\in\mathbb{R}^{3}$ and $\mathbf{v}_{t}\in\mathbb{R}^{3}$ define the robot’s position and velocity in the world frame, $R_{t}\in\mathbb{R}^{3\times 3}$ is the robot’s orientation matrix, and $\bm{\omega}_{t}\in\mathbb{R}^{3}$ is the angular velocity. To get the next state $\mathbf{s}_{t+1}$, in general, we need to solve the following ODE:

### v4

# FusionForce: End-to-end Differentiable Neural-Symbolic Layer for Trajectory Prediction

@@ Abstract
We propose end-to-end differentiable model that predicts robot trajectories on rough offroad terrain from camera images and/or lidar point clouds. The model integrates a learnable component that predicts robot-terrain interaction forces with a neural-symbolic layer that enforces the laws of classical mechanics and consequently improves generalization on out-of-distribution data. The neural-symbolic layer includes a differentiable physics engine that computes the robot’s trajectory by querying these forces at the points of contact with the terrain. As the proposed architecture comprises substantial geometrical and physics priors, the resulting model can also be seen as a learnable physics engine conditioned on real sensor data that delivers $10^{4}$ trajectories per second. We argue and empirically demonstrate that this architecture reduces the sim-to-real gap and mitigates out-of-distribution sensitivity. The differentiability, in conjunction with the rapid simulation speed, makes the model well-suited for various applications including model predictive control, trajectory shooting, supervised and reinforcement learning, or SLAM.
@@
Trajectory Prediction, Neural-Symbolic Models, Physics-based Learning, Off-road Navigation, Sensor Fusion
@@ I Introduction
Autonomous robotics in off-road environments holds immense promise for various applications, including outdoor logistics, inspections, and forestry operations. Yet, unlike systems designed for controlled environments, such as factories, off-road autonomous systems still remain prohibitively immature for dependable deployment. The main challenge that prevents reliable off-road deployment lies in the inability to predict the behavior of the robot on the terrain with a sufficient accuracy. Especially, the prediction of robot trajectory in tall vegetation or muddy and rocky terrains from camera images and lidar scans remains an unresolved problem in outdoor robotics. We address this issue by proposing an end-to-end differentiable physics-aware model that simultaneously predicts various terrain properties and potential robot trajectories, see Fig. 1 for an example.
Figure 1: Camera, predicted terrain properties and trajectories.
@@
Over the last decade, roboticists proposed a wide variety of white-box [1, 2, 3] and black-box models [4, 5, 6, 7, 8] for the off-road trajectory prediction. Black-box models primarily suffer from a severe out-of-distribution problem – the phenomenon where the distribution of training data does not correspond to the testing data. This problem naturally arises from the fact that the training data includes only the trajectories from safe terrain traversals, such as small terrain steps, while the decision about terrain traversability also naturally comprises the non-traversable terrains, such as tall cliffs, Fig. 3-top. Since the step-to-cliff-generalization of large state-of-the-art black-box models is typically poor, the safety of the resulting system is typically inferior. On the other hand, white-box architectures offer good generalization due to substantial inductive bias but often suffer from the sim-to-real gap – the phenomenon where a model """"trained in"""" or """"replaced by"""" simulation faces challenges or discrepancies when applied in the real world. Although several techniques, such as rapid motor adaptation [9], are known to partially suppress this issue through domain randomization, the sim-to-real gap, as well as the out-of-distribution problem, makes trajectory prediction from camera images prohibitively unreliable. We introduce a grey-box model that combines the best of both worlds to achieve better generalization and consequently better accuracy and smaller sim-to-real gap. This model enforces white-box laws of classical mechanics through a physics-aware neural symbolic layer, while preserving the ability to learn from huge data as it is end-to-end differentiable. We demonstrate that independently of the employed input sensor (camera, lidar or both), the proposed neural-symbolic physics layer consistently improves trajectory prediction error when compared to a state-of-the-art data-driven architecture such as LSTM, see Fig. 3 for an example.
[…]
In our implementation, the black-box part of our model predicts robot-terrain interaction forces and the true height at which they appear, while the neural symbolic layer, which contains a differentiable physics engine, solves the robot’s trajectory by querying these forces at the robot-terrain contacts. As the proposed architecture comprises substantial geometrical and physics priors, the resulting model can be also seen as learnable physics engine conditioned by a real image that delivers $10^{4}$ trajectories per second; see Fig. 2 for details.
@@
In addition to that, the model is end-to-end-differentiable; therefore, gradients can be backpropagated towards its (i) convolutional filters, (ii) camera and robot parameters, (iii) control, (iv) initial position, and (v) terrain properties. The differentiability, in conjunction with the rapid simulation speed, makes the model suitable for a myriad of tasks, including model predictive control [10], trajectory shooting [11], supervised and reinforcement learning [12], SLAM [13], online robot-model reidentification or camera recalibration [14]. The explainable structure of the proposed architecture also delivers a variety of intermediate outputs, such as terrain shape and its physical properties, robot-terrain reaction forces or contact points, which can all serve as efficient sources of self-supervision if measured during the training set creation or restricted in PINN-like manner [15].
In addition to that, we observed that the instability of gradient computation makes existing differentiable simulators such as NVIDIA’s WARP [16] and Google’s BRAX [17] prohibitively unreliable and slow for both learning and inference. As a suitable replacement, we propose our own differentiable neural symbolic layer, which integrates from-scratch implemented physics engine, into the proposed architecture. This solution outperforms existing works in gradient stability and computational speed while preserving sufficient accuracy for reliable trajectory prediction. As all target applications, including learning, control, planning, and SLAM, can be naturally parallelized, we also achieve significant speed-up through massive parallelization on GPU.
We emphasize that all state-of-the-art models for trajectory prediction of ground robots, including ours, are inherently statistically inconsistent because training data cannot, in principle, cover hazardous scenarios. As a result, the common practice of using a near-infinite number of training examples, as seen in large foundation models [18], does not guarantee superior performance. Addressing this issue without relying on real-world, robot-damaging trajectories remains an unresolved challenge. In this work, we propose a grey-box model that progresses toward addressing this issue by combining: (i) the strength of black-box models to learn efficiently from real-world data, thus reducing the sim-to-real gap, and (ii) the strengths of white-box approaches to generalize to unseen data, thereby addressing the out-of-distribution problem.
@@
Neural-symbolic physics-aware layer, which introduces a substantial physics prior that improves generalization on out-of-distribution terrains – a crucial property for any reliable deployment.
Explainability of the proposed grey-box model provides several well-interpretable intermediate outputs, such as contacts, slippage, or forces, that may serve as a natural source of self-supervision.
Image-conditioned simulation: The end-to-end differentiable image-conditioned simulation is suitable for a myriad of robotics tasks such as control, learning or SLAM, as it predicts a vast number of trajectories in parallel.
Experimental evaluation on non-rigid terrains: The proposed model outperforms other state-of-the-art methods especially on non-rigid terrains, such as grass or soft undergrowth that deforms when traversed by the robot. See Table I for quantitative results. All codes and data are publicly available [^1] https://github.com/ctu-vras/fusionforce.
Relation to Previous Work: A preliminary version of this work was presented at IROS 2024 [19]. This journal submission substantially extends the prior work by:
  • Proposing novel architecture that allows to seamlessly integrate 2D and 3D sensor modalities, such as camera and lidar. It was shown that including the lidar data improves the accuracy of the estimation of terrain properties by 16%, see Fig. 4 for an example.
  • Extending the predicted terrain properties by terrain friction, which significantly decreases the trajectory prediction error both in translation and in orientation (Table I). See Fig. 1 for an example of predicted terrain properties.
  • Improving the robot model for more accurate contact modeling and physical simulation. The new model is built from the robot’s URDF and uses 223 robot body points in comparison to only 16 points used in the approximate robot model from [19]. See also qualitative results in Fig. 11.
  • Benchmarking against a purely data-driven baseline. It is shown that the proposed physics engine outperforms the LSTM-based baseline up to 56% in translation and 48% in orientation in terms of trajectory prediction accuracy, Fig. 12.
Figure 4: Influence of lidar data Input images (first row). The predicted supporting terrain from lidar + camera (second and third row) - see 3D pointcloud to compare accuracy. The predicted supporting terrain from camera only (last row).
@@ II Related work
Recent work has shown that learning can yield extremely robust locomotion controllers despite complex terrain dynamics. For example, Lee et al. [20] train a neural controller for the ANYmal quadruped using only joint-encoder and IMU inputs, achieving robust zero-shot generalization to highly irregular outdoor terrain (mud, sand, vegetation, snow). Similarly, Choi et al. [21] incorporate a parameterized physics model of granular media into the reinforcement learning loop, enabling a quadruped to run on soft sand and generalize to diverse surfaces (asphalt, grass, synthetic mat) by identifying substrate parameters on the fly. These works focus only on proprioceptive feedback and do not allow to predict possibly fatal behaviour from exteroceptive sensors such as camera or LiDAR.
Complementary approaches use exteroceptive sensing to predict terrain traversability. Hoeller et al. [22] (ANYmal Parkour) employ onboard cameras and LiDAR to build a point-cloud map and latent terrain representation, guiding a high-level policy that selects from learned locomotion skills (walk, jump, climb, crouch) as the robot navigates obstacles. In wheeled off-road driving, Cai et al. [23] (EVORA) train a deep model to predict distributions of wheel traction from a semantic–elevation map fused from images and LiDAR, and incorporate this into a model-predictive planner that avoids low-traction or high-uncertainty regions. Vecchio et al. [24] use self-supervised learning on synthetic RGB video to regress per-region traversability costs, bridging sim-to-real with domain adaptation so the robot can estimate navigational cost maps from raw camera images. These perceptive methods produce rich terrain-cost models from sensor data, but typically treat the mapping from visuals to robot state as a black box, with limited explicit physics.
A growing trend is to extend learned models with physical reasoning. For instance, the deformable-terrain work by Choi et al. [21] embeds a physics model of sand within the controller, and EVORA explicitly models traction predictions to detect novel or out-of-distribution terrains. The AnyNav frameworkZ[25] exemplifies this neuro-symbolic direction: it uses a neural network to predict a friction coefficient from camera images that is grounded by a symbolic friction model, and then a friction-aware planner chooses safe, physically-feasible paths. Such approaches improve generalization by enforcing consistency with physical laws. In summary, while recent literature has significantly advanced data-driven traversability estimation and control, a gap remains in merging perception with explicit physics. The proposed physics-aware model aims to fill this gap by learning interpretable terrain parameters (e.g. friction, stiffness see Fig. 1 for an example) from sensors and plugging them into neural-symbolic layer that delivers expected trajectories by solving the underlying ODE, thereby enhancing robustness on out-of-distribution terrains. Since our primary focus is on developing a reliable model for trajectory prediction, we discuss related works mainly from the perspective of terrain and trajectory modeling, while details of their specific use for control implementations are omitted.
II-A Black-box (Data-driven) Models
The modeling of environment representation from single [26] or multiple RGB images [27] in a top-down view has been widely studied in the literature. The geometry of the scene prediction (taking into account occlusions) from stereo or depth-camera input has been addressed in [28]. The importance and benefits of sensor fusion (camera images, lidar, and radar) for environment representation have been demonstrated in [29]. The authors of RoadRunner [30] developed an end-to-end learning-based framework that predicts the traversability and elevation maps from multiple images and a lidar voxel map. However, the robot-terrain interaction and trajectory prediction tasks are not considered in these works.
The Recurrent Neural Networks (RNN) models [31] and especially their variant Long Short-Term Memory (LSTM) [32] and Gated Recurrent Unit (GRU) [33] have been widely and successfully used [34, 35, 36] for trajectory prediction tasks of various agents (mobile robots, cars, airplanes, people, and more). However, their main drawback lies in the difficulty of capturing the underlying physical laws, which leads to limited generalization capabilities
@@ II-C Grey-box (Hybrid) Models
III Methodology
A detailed overview of the proposed architecture that converts images, pointclouds and control commands into trajectories is depicted in Fig. 5. The model consists of several learnable modules that deeply interact with each other. The terrain encoder carefully transforms visual features from the input image into the heightmap domain using known camera geometry. These features are concatenated in the heightmap domain with voxel features estimated from lidar pointclouds. The resultant heightmap features are further refined by the multihead encoder-decoder architecture into interpretable physical quantities that capture properties of the terrain such as its shape, friction, stiffness, and damping. Next, the neuro-symbolic physics engine combines the terrain properties with the robot model, robot state, and control commands and delivers reaction forces at points of robot-terrain contacts. It then solves the equations of motion dynamics by integrating these forces and delivers the trajectory of the robot. Since the complete computational graph of the feedforward pass is retained, the backpropagation from an arbitrary loss, constructed on top of delivered trajectories, or any other intermediate outputs is at hand.
@@ III-A Terrain Encoder
The part of the FusionForce architecture (Fig. 5) that predicts terrain properties $\mathbf{m}$ from sensor measurements $\mathbf{z}$ is called terrain encoder. The proposed architecture starts by converting pixels from a 2D image plane into a heightmap with visual features. Since the camera is calibrated, we leverage this substantial geometrical prior to connect heightmap bins with pixels. We use known camera intrinsic parameters $\mathbf{K}\in\mathbb{R}^{3\times 3}$ to estimate rays corresponding to particular pixels – pixel-wise rays. For $i$-th ray, the convolutional network then predicts depth probabilities $p(d_{i})$ and visual features $\Phi_{i}$. Given the depth $d_{i}$ in pixel $(u_{i},v_{i})$, we reconstruct the 3D position $\mathbf{P}_{i}$ as follows:
$\mathbf{P}_{i}=d_{i}\cdot\mathbf{K}^{-1}\begin{bmatrix}u_{i}\\ v_{i}\\ 1\end{bmatrix},$ (1)
therefore all visual features for all nonzero depth probabilities $p(d_{i})>0$ can be vertically projected on a virtual heightmap. Finally, the $p(d_{i})$-weighted sum of visual features over each heightmap cell is computed. In particular, denoting $J$ the set of all 3D positions that are vertically projected into the heightmap cell $j$, the visual features $\varphi_{j}$ corresponding to this cell are computed as follows
$\varphi_{j}=\frac{\sum_{i\in J}\Phi_{i}p(d_{i})}{\sum_{i\in J}p(d_{i})}.$ (2)
Similarly, pointclouds are transformed into voxels and voxel-wise features are also vertically projected on heightmap domain. The resulting multichannel array that appears as the concatenation of camera and lidar features in heightmap domain, is further refined by deep convolutional network with encoder-decoder architecture to estimate the terrain properties $\mathbf{m}$ physics quantities that influences the robot motion.
The terrain properties include the geometrical heightmap $\mathcal{H}_{g}$, supporting terrain heightmap $\mathcal{H}_{t}$ (shape of the terrain layer hidden under the vegetation), terrain friction map $\mathcal{M}$, soft-terrain map $\Delta\mathcal{H}$ (models a partially flexible layer of terrain (e.g. mud) that is hidden under flexible vegetation), stiffness map $\mathcal{K}$ (strength of the terrain to prevent penetration by the robot), and dampening map $\mathcal{D}$ (the last two are considered constant in the current implementation).
III-B Differentiable Neuro-Symbolic Physics Engine
The differentiable physics engine solves the robot motion equation and estimates the trajectory corresponding to the delivered forces. The trajectory is defined as a sequence of robot states $\tau=\{s_{0},s_{1},\ldots,s_{T}\}$, where $\mathbf{s}_{t}=[\mathbf{x}_{t},\mathbf{v}_{t},R_{t},\bm{\omega}_{t}]$ is the robot state at time $t$, $\mathbf{x}_{t}\in\mathbb{R}^{3}$ and $\mathbf{v}_{t}\in\mathbb{R}^{3}$ define the robot’s position and linear velocity in the world frame, $R_{t}\in\mathbb{R}^{3\times 3}$ is the robot’s orientation matrix, and $\bm{\omega}_{t}\in\mathbb{R}^{3}$ is the angular velocity. To get the next state $\mathbf{s}_{t+1}$, in general, we need to solve the following ODE: